Wednesday, August 30, 2006

Obtaining the Xe(L) spetra successfully

The laser system ran very well today. The front end output a laser beam with about 26 mJ energy. After the Prometheus amplifier, the energy reached about 360 mJ as energizing 25 kV high voltage. Today the total shots is 1218, no misfire happened. The filament was shut down in the beginning, I could hear the click from the filament relay, after turn on again, the problem disappeared.

Because the nozzle was changed, I adjusted the timing delay for optimizing the Xe(L) outputs. Alex took many spectra from the target side.

No comments: